OpenStudy (anonymous):

Ginny's class ordered 2 pizzas. They ate seven-eighths of each pizza. How many pizzas did they eat?

OpenStudy (ganpat):

7/8 * 2 = 7/4 = 1 3/4 = 1 and quarter pizza..

OpenStudy (anonymous):

thx

OpenStudy (ganpat):

w/c

OpenStudy (anonymous):

can someone help me with a question im new